# DocHPLT: A Massively Multilingual Document-Level Translation Dataset

### v1

@@ Abstract
Existing document-level machine translation resources are only available for a handful of languages, mostly high-resourced ones. To facilitate the training and evaluation of document-level translation and, more broadly, long-context modeling for global communities, we create DocHPLT, the largest publicly available document-level translation dataset to date. It contains 124 million aligned document pairs across 50 languages paired with English, comprising 4.26 billion sentences, with further possibility to provide 2500 bonus pairs not involving English. Unlike previous reconstruction-based approaches that piece together documents from sentence-level data, we modify an existing web extraction pipeline to preserve complete document integrity from the source, retaining all content including unaligned portions. After our preliminary experiments identify the optimal training context strategy for document-level translation, we demonstrate that LLMs fine-tuned on DocHPLT substantially outperform off-the-shelf instruction-tuned baselines, with particularly dramatic improvements for under-resourced languages. We open-source the dataset under a permissive license, providing essential infrastructure for advancing multilingual document-level translation.
@@ 1 Introduction
The field of natural language processing (NLP) is shifting its focus toward end-to-end, complex tasks. This increases the demand for techniques and resources beyond the sentence level, with document-level machine translation (DocMT) being a prime example (Maruf and Haffari, 2018; Zhang et al., 2018; Agrawal et al., 2018; Huo et al., 2020). DocMT requires models to translate an entire document as a coherent unit rather than isolated segments. This approach is necessary for handling various discourse phenomena: anaphora, deixis, ellipsis, discourse connectives, grammatical and lexical cohesion (Maruf et al., 2021), which sentence-level translation typically loses (Müller et al., 2018; Bawden et al., 2018; Voita et al., 2018). Recent long-context large language models (LLMs) are well-suited for this task, as they are usually pre-trained to process thousands of tokens. However, DocMT remains largely unexplored or untested for most languages due to a simple but significant problem: we lack document-level parallel data for training and evaluation.
@@
Unlike sentence pairs that can be easily harvested from web sources, intact document translations are rarer in digital formats suitable for automatic extraction than their sentence-level counterparts. Manual creation through professional translation is prohibitively expensive for most language pairs. While a handful of language pairs have some document-level MT resources, the majority of languages have none. This creates two related problems at once: we cannot build DocMT for these languages, and we cannot evaluate DocMT properly. As NLP research moves toward more end-to-end, context-aware applications, this data gap means that most languages get left behind.
To address this problem, we introduce DocHPLT, a large multilingual document-level translation dataset covering 50 language pairs with English. The resulting corpus contains 87.8 million documents across 50 languages—a total of 4.3 billion sentences. One highlight of our work is the focus on including medium- and low-resource languages that previous DocMT datasets have overlooked. Practitioners can also use English as a pivot to obtain up to 2500 extra non-English-centric language pairs, expanding the dataset’s usefulness beyond English-centric translation.
@@
Our methodology differs from the majority of previous efforts that reconstruct documents from sentence pairs after the fact. Instead, we modify the web extraction pipeline itself to preserve document structure from the beginning, retaining documents in their entirety with all original context and non-parallel text. For each language pair, we deliver the aligned documents along with quality scores at the sentence level, alignment density metrics, and the original document structure.
@@
Using DocHPLT, we conduct extensive experiments with different modeling methods for document-level translation. We first try different context sizes for LLM fine-tuning: 1) full document-to-document training with loss calculated on the entire target; and 2) chunk-based training with loss computed on individual segments. These experiments determine the optimal context granularity for training DocMT. Then, in addition to prompting off-the-shelf instruction-tuned large language models (LLMs) as a baseline, we run monolingual and multilingual fine-tuning using DocHPLT. The usefulness of our data is reflected empirically. LLMs fine-tuned on our data consistently outperform prompting baselines, proving that, with our data, practitioners gain strong performance in document-level translation for languages often considered “unsupported” in the machine translation research community.
In summary, our contributions, centred around DocHPLT, are as follows:
@@
  • Scale and diversity: DocHPLT is the largest publicly available document-level translation resource: 124M document pairs for 50 languages, totaling 4.26B sentences, with extensive medium- and low-resource coverage.
@@
  • Document-first approach: Instead of piecing together documents from sentence pairs, we preserve complete documents with original structure and unaligned text, enriched with quality metrics such as alignment density and sentence pair-level scores.
@@
  • Empirical validation: Through LLM experiments on both the internal test set and WMT24++, we establish baselines, test various training strategies, and demonstrate gains in DocMT using our data.
@@ 2.1 Document-Level Translation
Document-level translation aims to process an entire document as a coherent unit, rather than processing each sentence independently. This paradigm leverages the ability of modern neural architectures, lately LLMs, to handle long context, making it particularly effective for capturing document-level discourse structures. Although there have been several massive-scale parallel corpus mining efforts (Bañón et al., 2020; El-Kishky et al., 2020; Schwenk et al., 2021; Arefyev et al., 2024), document-level parallel data remains scarce.
Recent work has demonstrated that going beyond sentence-level translation is essential for handling discourse phenomena such as coreference resolution (Müller et al., 2018; Bawden et al., 2018; Voita et al., 2018). The development of dedicated document-level benchmarks further reflects this growing interest in evaluating MT systems in context (Guillou and Hardmeier, 2016; Jwalapuram et al., 2020; Wicks and Post, 2023; Fernandes et al., 2023).
@@
2.2 Document-Level Data Provision
Existing document-level datasets remain limited in size and scope, particularly when extending beyond English-centric or high-resource languages. Moreover, there is little agreement on what constitutes a “document”; definitions vary widely across studies, ranging from short paragraphs to entire articles or books. This lack of standardization, combined with the scarcity of large-scale multilingual document-level corpora, motivates the need for more diverse resources such as the one we present in this work. Before illustrating our data methodology, we survey two typical methods used in creating document-level translation data.
@@ 2.2 Document-Level Data Provision
A common strategy to obtain document-level parallel data is to reconstruct from existing sentence-level corpora. Notably, the sentence-level ParaCrawl (Bañón et al., 2020) has been widely used as a seed for this purpose. Al Ghussin et al. (2023) extracted English–German parallel paragraphs from ParaCrawl, although these are not full-document units. ParaDocs (Wicks et al., 2024) recovered document-level data from ParaCrawl, News Commentary (Kocmi et al., 2023), and Europarl (Koehn, 2005) for German, French, Spanish, Italian, Polish, and Portuguese, all paired with English. Similarly, Pal et al. (2024) released a large-scale reconstructed corpus for German, French, Czech, Polish, and Russian—again all paired with English, along with an open-source pipeline for extension to other languages. Recently, Ramos et al. (2025) presented DocBlocks, a multilingual document-level dataset covering several high-resource languages by aggregating and applying a rigorous cleaning pipeline to data from public sources.
@@
An alternative approach is to collect or create document-level parallel corpora directly from scratch. Literary works have been a popular origin. Jiang et al. (2022) introduced a Chinese–English corpus based on web novels, where each chapter, with a median of 30 sentences, is treated as a document. Thai et al. (2022) created PAR3 by aligning machine and human translations of 118 novels across 19 languages at the paragraph level. Jin et al. (2024) constructed JAM from 160 English-Chinese novel pairs with chapter-level alignment. More recently, Alabi et al. (2025) created AFRIDOC-MT, a document-level translation corpus sourced from IT news and health articles and manually translated from English. By covering Amharic, Hausa, Swahili, Yorùbá, and Zulu, it extends DocMT data to medium and lower-resourced languages. Such data, directly derived from resources intended to be document-aligned, is high-quality but often limited by languages due to the cost and effort required.
@@
Our work gathers document translations from large web crawls, but differs fundamentally from reconstruction-based approaches. As explained later in Section 3, rather than piecing together documents from sentence pairs post-hoc, we modify the document alignment stage of the extraction pipeline to preserve complete document structure from the beginning. This document-first methodology ensures we retain all original content including unaligned portions, positioning our approach as collection-based at the document level while leveraging existing text crawling and processing infrastructure.
@@ 3 DocHPLT
In this section, we explain how we modify and then apply an existing parallel sentence extraction pipeline from ParaCrawl Bañón et al. (2020) to extract a document-level corpus from a large multilingual web crawl, HPLT (Burchell et al., 2025).
@@ 3.1 Dataset Creation
The starting point for our dataset creation is 15TB of cleaned web documents derived from the Internet Archive [^1] https://archive.org/ and CommonCrawl [^2] https://commoncrawl.org/ as an outcome of the HPLT corpus (Burchell et al., 2025). In the preparation of HPLT, the document text was extracted from HTML using Trafilatura (Barbaresi, 2021) and language-classified using openLID (Burchell et al., 2023). In this work, a document is defined as the full text content retrieved from archive snapshots of a specific URL.
To extract a parallel corpus of documents, we use a modified version of the ParaCrawl extraction pipeline (Bañón et al., 2020). The original pipeline is sentence-oriented, i.e. it produces a sentence-aligned corpus and discards unaligned sentences. Yet, the pipeline runs document alignment followed by sentence alignment in separate stages, allowing us to intervene to produce document-oriented data. We extract and record each pair of aligned documents, then map sentence alignment back to this complete document structure. This document-first methodology ensures we retain all document content, even unaligned portions, to provide richer context than traditional parallel corpora.
@@
We transform each document into a hierarchical XML representation that preserves its complete internal structure. Paragraphs are split by newline characters and maintain their original boundaries, while the text of each paragraph is segmented into sentences using the Loomchild Segmenter (Miłkowski and Lipski, 2011). Every structural element receives a unique identifier with paragraphs, such as <P id="4">, and sentences, such as <s id="4.3">. This structured representation allows us to track alignments at both document and sentence levels while retaining all content from the original HPLT documents.
@@
Since our initial collection contains multiple temporal snapshots of the same URLs, we implement a content-based deduplication strategy. First, within each language-specific collection, we remove duplicates by grouping documents by their source URL and then comparing the exact-match content. This ensures we keep only unique document versions for each URL. Second, we perform global deduplication across all English documents from the 50 language pairs, consolidating them into a single collection. This is necessary because the same English document may appear in multiple language pairs (e.g., the same English page aligned to both Basque and Catalan translations). After deduplication, we have a clean collection of unique documents for each of the 50 source languages and a single, unified collection for all English documents, ensuring each unique document version appears exactly once while preserving all alignment relationships.
We deliberately preserve duplicate content across different URLs and retain near-duplicates within the same URL. This design choice maximizes researcher flexibility by allowing downstream users to apply filtering strategies suited to their particular use cases. Additionally, duplicate content from different URLs preserves valuable metadata, particularly the source URL itself, which may indicate different domains, publication contexts, or content distribution patterns. Near-duplicates may represent meaningful content variations such as updates, revisions, or editorial differences that could be relevant for studying linguistic evolution or content dynamics.
@@
The ParaCrawl pipeline originally used MinHash (Broder, 1997) to deduplicate similar sentences, grouping them together and assigning the same quality scores regardless of their source documents. We modify this step to remove MinHash deduplication entirely, instead maintaining all original texts and tracking which documents they came from. This allows us to preserve the complete document structure while still computing alignment quality scores—BLEUalign (Sennrich and Volk, 2010), Bicleaner, and Bifixer (Ramírez-Sánchez et al., 2020)—for each sentence pair. Figure 1 illustrates examples of good, bad, and multi-way alignments along with their corresponding quality scores. We then map each alignment back to its specific source and target documents, maintaining the document-sentence relationships throughout the process. For any document that had multiple versions of the same URL, we explicitly check that every sentence referenced in an alignment link actually exists in the final XML file to ensure it references the correct version(s). The output follows the standard cesAlign XML format [^3] https://opus.nlpl.eu/legacy/trac/wiki/DataFormats.html, where each alignment links specific sentence IDs between source and target documents along with their quality scores.
@@
As a “bonus” data release, English can be used as a pivot language to derive a corpus beyond the English-centric pairs. This enables the modeling and evaluation of DocMT between two non-English languages. The process is straightforward: if a document in a language and another document in another language are both aligned to the same English document, there is a direct alignment between the two documents.
@@ 3.2 Data Statistics
Across all language pairs, DocHPLT contains 87.8 million unique documents with 4.3 billion total sentences, averaging 48.6 sentences per document. The English collection dominates with 47.5 million documents (2.67 billion sentences), while individual non-English languages range from Japanese with 4 million documents (164 million sentences) down to Xhosa with 22 thousand documents (996 thousand sentences). These documents form 124 million aligned document pairs, with an average of 14.8 sentence-level alignments per document pair. Document coverage varies significantly: Japanese-English and Turkish-English each contribute over 11 million aligned document pairs, respectively, while under-represented languages like Sinhala-English (123 thousand document pairs), Uzbek-English (157 thousand document pairs), and Xhosa-English (44 thousand document pairs) have substantially smaller collections.
We observe considerable variations in document length ratios between aligned pairs. Malayalam-English has the highest ratio (3.91), indicating substantially longer English documents, while Arabic-English shows a lower ratio (0.84), suggesting shorter English documents relative to Arabic. Additionally, the average Bicleaner scores vary significantly, with language pairs like Arabic-English (0.700) demonstrating relatively high-quality alignments, whereas pairs such as Maltese-English (0.293) display substantially lower average alignment quality.
@@
This feature may reflect the quality and the characteristics of the documents: a very high AD might suggest that the documents were machine-translated (at the sentence level), whereas a very low AD might imply that they were accidentally matched, possibly due to high-frequency phrases or placeholders. We observe considerable variation in AD across language pairs, e.g., Welsh-English (cy-en) and Afrikaans-English (af-en) show notably high average alignment densities (0.426 and 0.446, respectively), compared to languages like Farsi, Malayalam, and Marathi, where alignments are much sparser (0.153, 0.151, and 0.150, respectively).

### v2

@@ Abstract
Existing document-level machine translation resources are only available for a handful of languages, mostly high-resourced ones. To facilitate the training and evaluation of document-level translation and, more broadly, long-context modeling for global communities, we create DocHPLT, the largest publicly available document-level translation dataset to date. It contains 124 million aligned document pairs across 50 languages paired with English, comprising 4.26 billion sentences. By adding pivoted alignments, practitioners can obtain 2500 additional pairs not involving English. Unlike previous reconstruction-based approaches that piece together documents from sentence-level data, we modify an existing web extraction pipeline to preserve complete document integrity from the source, retaining all content, including unaligned portions. After our preliminary experiments identify the optimal training context strategy for document-level translation, we demonstrate that LLMs fine-tuned on DocHPLT substantially outperform off-the-shelf instruction-tuned baselines, with particularly dramatic improvements for under-resourced languages. We open-source the dataset under a permissive license, providing essential infrastructure for advancing multilingual document-level translation.
@@ 1 Introduction
The field of natural language processing (NLP) is shifting its focus toward end-to-end, complex tasks, including the domain of machine translation. This increases the demand for techniques and resources beyond the sentence level, with document-level machine translation (DocMT) being a prime example (Maruf and Haffari, 2018; Zhang et al., 2018; Agrawal et al., 2018; Huo et al., 2020). While there is not a single definition of a document, DocMT requires models to translate more than one sentence as a coherent unit rather than isolated segments. This approach is necessary for handling various discourse phenomena: anaphora, deixis, ellipsis, discourse connectives, grammatical and lexical cohesion (Maruf et al., 2021), which sentence-level translation typically loses (Müller et al., 2018; Bawden et al., 2018; Voita et al., 2018). Recent long-context large language models (LLMs) are well-suited for this task, as they are usually pre-trained to process thousands of tokens at a time. However, DocMT remains largely unexplored or untested for most languages due to a simple but significant problem: we lack document-level parallel data for both model building and evaluation.
@@
Historically, parallel corpora were mostly constructed in a sentence-oriented manner, using a pipeline that split the text into sentences, aligned them, and then discarded unaligned and multiply-aligned sentences. While a handful of language pairs have some document-level MT resources, the majority of languages have none. This creates two related problems at once: we cannot build DocMT for these languages, and we cannot evaluate DocMT properly. As NLP research moves toward more end-to-end, context-aware applications, this data gap means that most languages get left behind.
@@
We tackle this problem by extracting parallel documents from large web crawls, but our methodology differs from the majority of previous efforts that reconstruct data from sentence pairs after the fact. Instead, we modify the web extraction pipeline itself to preserve document structure from the beginning, retaining documents in their entirety with all original context and non-parallel text. For each language pair, we deliver the aligned documents along with quality-scored sentence alignments and alignment density metrics.
Our effort yielded DocHPLT, a large multilingual document-level translation dataset covering 50 language pairs with English, listed in Appendix A. The resulting corpus contains 87.8 million documents in English and 50 other languages, 124 million aligned document pairs, and 4.26 billion sentences. A highlight of our work is the focus on medium- and low-resource languages that previous DocMT datasets have overlooked. Practitioners can also use English as a pivot to align up to 2500 extra non-English pairs, expanding the dataset’s usefulness beyond English-centric translation.
@@
Using DocHPLT, we conduct extensive experiments with different modeling methods for LLM-based document-level translation. We first try different context sizes for LLM fine-tuning: 1) full document-to-document training with loss calculated on the entire target; and 2) chunk-based training with loss computed on individual segments. These experiments determine the optimal context granularity for our subsequent work. Then, in addition to prompting off-the-shelf instruction-tuned large language models (LLMs) as a baseline, we run monolingual and multilingual fine-tuning using DocHPLT, tested on both seen and unseen languages. The usefulness of our data is reflected empirically: LLMs fine-tuned on our data consistently outperform prompting baselines, showing that practitioners can gain strong performance in DocMT for languages often considered “unsupported” in the machine translation research community.
In summary, our contributions, centred around the DocHPLT resource, are as follows:
@@
  • Scale and diversity: DocHPLT is the largest publicly available document-level translation resource: 124M document pairs for 50 languages paired with English, totaling 4.26B sentences, with extensive medium- and low-resource coverage.
@@
  • Document-first approach: Instead of piecing together documents from aligned sentence pairs, we preserve complete documents with original structure and unaligned text, enriched with quality metrics such as alignment density and sentence pair-level scores.
@@
  • Empirical validation: Through LLM experiments on both the internal test set and WMT24++, we establish baselines, test different training strategies, and demonstrate gains in DocMT using our data.
@@ 2.1 Document-Level Translation
Document-level translation aims to process an entire document as a coherent unit, rather than processing each sentence independently. This paradigm leverages the ability of modern neural architectures, lately LLMs, to handle long context, making it particularly effective for capturing document-level discourse structures. Recent work has demonstrated that going beyond sentence-level translation is essential for handling discourse phenomena such as coreference resolution (Müller et al., 2018; Bawden et al., 2018; Voita et al., 2018). The development of dedicated document-level benchmarks further reflects this growing interest in evaluating MT systems in context (Guillou and Hardmeier, 2016; Jwalapuram et al., 2020; Wicks and Post, 2023; Fernandes et al., 2023).
@@
2.2 Document-Level Translation Data
Although there have been several massive-scale parallel corpus mining efforts (Bañón et al., 2020; El-Kishky et al., 2020; Schwenk et al., 2021; de Gibert et al., 2024; Burchell et al., 2025), document-level data remain limited in size and scope, particularly when extending beyond English-centric or high-resource languages. Moreover, there is little agreement on what constitutes a “document”; definitions vary widely across studies, ranging from short paragraphs to entire articles or books. This lack of standardization, combined with the scarcity of large-scale multilingual document-level corpora, motivates the need for more diverse resources such as the one we present in this work. Before illustrating our data methodology, we survey two typical methods used in creating document-level translation data.
@@ 2.2 Document-Level Translation Data
A common strategy to obtain document-level parallel data is to reconstruct from existing sentence-level data. Notably, the sentence-level ParaCrawl (Bañón et al., 2020) has been widely used as a seed for this purpose. Al Ghussin et al. (2023) extracted English–German parallel paragraphs from ParaCrawl, although these are not full-document units. ParaDocs (Wicks et al., 2024) recovered document-level data from ParaCrawl, News Commentary (Kocmi et al., 2023), and Europarl (Koehn, 2005) for German, French, Spanish, Italian, Polish, and Portuguese, all paired with English. Similarly, Pal et al. (2024) released a large-scale reconstructed corpus for German, French, Czech, Polish, and Russian—again all paired with English, along with an open-source pipeline for extension to other languages.
@@
An alternative approach is to collect or create document-level parallel corpora directly from targeted sources from scratch. Earlier efforts include Europarl based on the proceedings of the European Parliament (Koehn, 2005) and OpenSubtitles from movie and TV subtitles (Lison and Tiedemann, 2016), but these essentially consist of “spoken” documents, where the former is divided into speeches and the latter into films/shows. Literary works have also been a popular origin. Jiang et al. (2022) introduced a Chinese–English corpus based on web novels, where each chapter, with a median of 30 sentences, is treated as a document. Thai et al. (2022) created PAR3 by aligning machine and human translations of 118 novels across 19 languages at the paragraph level. Jin et al. (2024) constructed JAM from 160 English-Chinese novel pairs with chapter-level alignment. More recently, Alabi et al. (2025) created AFRIDOC-MT, a document-level translation corpus sourced from IT news and health articles and manually translated from English. By covering Amharic, Hausa, Swahili, Yorùbá, and Zulu, it extends DocMT data to medium and lower-resourced languages. Wastl et al. (2025) scraped a Swiss online news outlet to create 20min-XD, a French-German dataset. Such data, directly derived from resources intended to be document-aligned, is high-quality but often limited by languages due to the coverage of the upstream source and/or the cost and effort required.
@@
Our work gathers document translations from large web crawls, but differs fundamentally from reconstruction-based approaches. As explained later in Section 3, rather than piecing together documents from sentence pairs post-hoc, we modify the document alignment stage of the extraction pipeline to preserve complete document structure from the beginning. This document-first methodology ensures we retain all original content, including unaligned portions, positioning our approach as collection-based at the document level while leveraging existing text crawling and processing infrastructure.
@@ 3 DocHPLT
In this section, we explain how we modify and then apply an existing parallel sentence extraction pipeline from ParaCrawl to extract a document-level corpus from a large multilingual web crawl, HPLT.
@@ 3.1 Dataset Creation
The starting point for our dataset creation is 15TB of cleaned web documents derived from the Internet Archive [^1] https://archive.org/ and CommonCrawl [^2] https://commoncrawl.org/ released as version 2 of the HPLT corpus (Burchell et al., 2025). In the preparation of HPLT, the document text was extracted from HTML using Trafilatura (Barbaresi, 2021) and language-classified using openLID (Burchell et al., 2023). In this work, a document is defined as the full text content retrieved from archive snapshots of a specific URL.
To extract a parallel corpus of documents, we use a modified version of the ParaCrawl extraction pipeline (Bañón et al., 2020). The original pipeline is sentence-oriented, i.e. it produces a sentence-aligned corpus and discards unaligned sentences. But because the pipeline runs document alignment followed by sentence alignment in separate stages, we are able to intervene to produce document-oriented data. We extract and record each pair of aligned documents, then map the unfiltered sentence alignments back into their source documents. This document-first methodology ensures we retain all document content, even unaligned portions, to provide richer context than traditional parallel corpora.
@@
We transform each document into a hierarchical XML representation that preserves its internal structure. Paragraphs are split by newline characters and maintain their original boundaries, while the text of each paragraph is segmented into sentences using the Loomchild Segmenter (Miłkowski and Lipski, 2011). Every structural element receives a unique identifier with paragraphs, such as <P id="4">, and sentences, such as <s id="4.3">. This structured representation allows us to track alignments at both document and sentence levels while retaining all content from the original HPLT documents.
@@
Since our initial collection contains multiple temporal snapshots of the same URLs, we implement a content-based deduplication strategy. First, within each language-specific collection, we remove duplicates, using the URL together with the full text as the key. This ensures we keep only unique document versions for each URL. Second, we perform global deduplication, based on the URL and text as the key, across all English documents from the 50 language pairs, consolidating them into a single collection. This is necessary because the same English document may appear in multiple language pairs (e.g., the same English page aligned to both Basque and Catalan translations). After deduplication, we have a clean collection of unique documents for each of the 50 source languages and a single, unified collection for all English documents, ensuring each unique document version appears exactly once while preserving all alignment relationships.
We deliberately preserve duplicate content across different URLs and retain near-duplicates within the same URL. This design choice maximizes research flexibility by allowing downstream users to apply filtering strategies suited to their particular use cases. Additionally, duplicate content from different URLs preserves valuable metadata, particularly the source URL, which may indicate different domains, publication contexts, or content distribution patterns. Near-duplicates also represent meaningful content variations such as updates, revisions, or editorial differences. Deduplicating content only for each language separately results in a 3.3% drop in our document count from the original DocHPLT corpus (see Appendix A).
@@
The ParaCrawl pipeline originally used MinHash (Broder, 1997) to deduplicate similar sentences, grouping them and assigning the same quality scores regardless of their source documents. We modify this step to remove MinHash deduplication entirely, instead maintaining all original texts and tracking which documents they came from. This allows us to preserve the complete document structure while still computing alignment quality scores—BLEUalign (Sennrich and Volk, 2010), Bicleaner, and Bifixer (Ramírez-Sánchez et al., 2020)—for each sentence pair. The document in Figure 1 illustrates examples of good, bad, and multi-way alignments along with their corresponding quality scores. We then map each alignment back to its specific source and target documents, maintaining the document-sentence relationships throughout the process. For any document that had multiple versions with the same URL, we explicitly check that every sentence referenced in an alignment link actually exists in the final XML file to ensure it references the correct version(s). The output follows the standard cesAlign XML format [^3] https://opus.nlpl.eu/legacy/trac/wiki/DataFormats.html, where each alignment links specific sentence IDs between source and target documents along with their quality scores.
@@
As a “bonus” data release, English can be used as a pivot language to derive a corpus beyond the English-centric pairs. This enables the modeling and evaluation of DocMT between two non-English languages. The process is straightforward: if a document in a language and another document in another language are both aligned to the same English document, then we assume a direct alignment between the two documents. We pivot the sentence alignments in a similar way.
@@ 3.2 Data Statistics
Across all language pairs, DocHPLT contains 87.8 million unique documents with 4.26 billion total sentences, averaging 48.6 sentences per document. The English collection dominates with 47.5 million documents (2.67 billion sentences), while individual non-English languages range from Japanese with 4 million documents (164 million sentences) down to Xhosa with 22 thousand documents (996 thousand sentences). These documents form 124 million aligned document pairs, with an average of 14.8 sentence-level alignments per document pair. Document coverage varies significantly: Japanese-English and Turkish-English each contribute over 11 million aligned document pairs, respectively, while under-represented languages like Sinhala-English (123 thousand document pairs), Uzbek-English (157 thousand document pairs), and Xhosa-English (44 thousand document pairs) have substantially smaller collections.
We observe considerable variations in document length ratios between aligned pairs, ranging from 3.91 (Malayalam-English), where the English documents are typically longer, to 0.84 (Arabic-English). Additionally, the average Bicleaner scores vary significantly, with language pairs like Arabic-English (0.700) demonstrating relatively high-quality alignments, whereas pairs such as Maltese-English (0.293) display substantially lower average alignment quality.
@@
Alignment density ranges between 0 (no aligned pairs) and 1 (perfect sentence-level coverage) if alignments are strictly one-to-one; however, since our alignment procedure allows one-to-many and many-to-one mappings, values above 1 are also possible. This feature may reveal the quality and the characteristics of the documents: an AD of exactly 1 could suggest that the documents were machine-translated (at the sentence level), whereas a very low AD might imply that they were accidentally matched, possibly due to high-frequency phrases or placeholders.
We observe considerable variation in AD across language pairs, e.g., Welsh-English (cy-en) and Afrikaans-English (af-en) show notably high average alignment densities (0.426 and 0.446, respectively), compared to languages like Farsi, Malayalam, and Marathi, where alignments are much sparser (0.153, 0.151, and 0.150, respectively). While some language pairs exhibit higher or lower densities, these scores are better understood relative to other scores rather than absolute terms.
We did not observe any consistent correlation between automatic quality metrics (BicleanerAI and CometKiwi) and AD values. Future work should investigate more carefully how AD should be interpreted and framed.